OpenStudy (anonymous):

Find the next 3 terms 5,12,26

OpenStudy (inkyvoyd):

It could be anything.

OpenStudy (inkyvoyd):

Please tell us what math you are learning right now.

OpenStudy (anonymous):

looks like 2x+2 so maybe \(2\times 26+2=54\) for the fist one

OpenStudy (inkyvoyd):

(a polynomial of fourth degree could make it any algebraic number)

OpenStudy (anonymous):

then \(2\times 54+2=110\) for the second

OpenStudy (anonymous):

satellite is right so 54, 110, 222

OpenStudy (inkyvoyd):

(a polynomial of fourth degree with transcendental or algebraic numbers as coefficients could make it anything...)

OpenStudy (anonymous):

@Numbers i wouldn't steer you wrong

OpenStudy (inkyvoyd):

it could be pi. it could be e. it could be i. it could be e^(ipi)

OpenStudy (anonymous):

54, 110, 222

OpenStudy (anonymous):

I agree with satellite :)

OpenStudy (inkyvoyd):

It looks like it... but who knows? (it probably is it)

OpenStudy (inkyvoyd):

a polynomial of seventh degree could make the numbers, e, i, pi.

OpenStudy (inkyvoyd):

So, I'm going with e, i, pi (although @satellite73 has what you are really looking for)